# Semantic Segmentation Algorithm Based on Light Field and LiDAR Fusion

### v1

@@ IV-B Ablation
TABLE II presents our ablation study results. The baseline extends Mseg3D by incorporating image branch annotation. Our alignment loss enhances image branch performance but slightly degrades the point cloud branch, likely due to partial misalignment caused by occlusion. Despite this, the overall mIoU improves by 0.71. Interpolating the point cloud feature map on the image plane significantly boosts image segmentation, indirectly benefiting the point cloud branch, leading to a 1.96 mIoU increase. Adding the depth difference perception module further enhances occluded object segmentation, raising mIoU by 2.71. Finally, integrating multi-view light field images improves the mIoU by another 2.71. These results validate the effectiveness of the point-pixel interpolation, depth difference perception, and light field integration.
@@ V CONCLUSION
In this work, we introduce the first multimodal dataset for real-world traffic scenes that integrates light field and LiDAR modalities. It includes 623×9 annotated light field images from a 3x3 camera array and corresponding LiDAR point cloud data across 623 frames. Unlike existing datasets, our camera array provides a broader perspective disparity, improving semantic segmentation accuracy. A key feature is the per-pixel annotation across all viewpoints, enabling comprehensive light field segmentation. We benchmark leading single-image, light-field and point-cloud segmentation techniques, showcasing the effectiveness of our dataset. Additionally, we introduce Mipfseg and Mlpfseg, the first multimodal fusion segmentation network for both point clouds and light field images. The network uses point-pixel interpolation fusion to create dense features from sparse point cloud data, with enhanced fusion via alignment loss and cross-attention mechanisms. The depth difference perception module strengthens the learning of occluded object parts, improving recognition accuracy. By incorporating light field images, our approach improves segmentation accuracy for small and occluded objects, achieving state-of-the-art performance for both image and point cloud modalities. In future work, we aim to integrate light field depth estimation for joint training, further refining the depth perception module and developing a more unified network.

### v2

@@ IV-B Ablation
Table III presents the ablation study results. The baseline extends MSeg3D by incorporating the image branch. Introducing the alignment loss improves image segmentation but slightly affects point cloud performance due to partial cross-modal misalignment, yielding an overall gain of 0.71 mIoU.
The proposed interpolation-based feature interaction significantly improves image segmentation and also benefits point cloud performance, leading to a 1.96 mIoU increase. Incorporating the depth difference perception module further enhances occluded object modeling, resulting in a 2.71 mIoU gain. Finally, integrating multi-view light-field images achieves the best performance with an additional improvement of 3.02 mIoU.
Overall, the results validate the effectiveness of the proposed interpolation, depth perception, and multi-view fusion components.
@@ V CONCLUSION
[…]